# CoCoA: A Generalized Approach to Uncertainty Quantification by Integrating Confidence and Consistency of LLM Outputs

### v2

@@ Abstract
Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompasses a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling. Several recent methods have combined these two approaches and shown impressive performance in various applications. However, they sometimes fail to outperform much simpler baseline methods. Our investigation reveals distinctive characteristics of LLMs as probabilistic models, which help to explain why these UQ methods underperform in certain tasks. Based on these findings, we propose a new way of synthesizing model confidence and output consistency that leads to a family of efficient and robust UQ methods. We evaluate our approach across a variety of tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advancements in information retrieval, question answering, machine translation, and other language-driven applications. As these models become an integral part of everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors or uncertainty can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which measures how confident a model is in its outputs, and makes possible the rejection of generations with a high risk of being incorrect.
UQ for LLMs is a rapidly advancing research area, with new methods for estimating uncertainty emerging each year. The large portion of novel techniques is based on two fundamental approaches: information-theoretic analysis and the assessment of output consistency.
@@
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020). A key limitation of these methods is that they cannot account for the semantic variability across multiple possible outputs for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationships between the sampled outputs (Lin et al., 2024; Fomicheva et al., 2020). Information-theoretic and consistency-based methods have complementary strengths: the former provides insights into the model’s internal confidence, while the latter captures the uncertainty as objective variability of meaning among sampled outputs. For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024). Although such methods show good performance in various applications, they sometimes fail to outperform their simpler counterparts in certain scenarios (Vashurin et al., 2024).
Our investigation revealed distinctive characteristics of LLMs as probabilistic models, shedding light on why current state-of-the-art UQ methods that attempt to integrate both approaches often underperform in certain tasks. Specifically, we highlight the complexity of the token prediction process and the absence of a unified framework that simultaneously addresses model confidence and output variability, both of which can limit the effectiveness of existing UQ techniques. This insight drives our proposal for a novel family of methods that integrate model confidence with output consistency, resulting in more efficient and robust UQ techniques. Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Our main contributions can be summarized as follows.
  • We identify key limitations in current UQ methods for LLMs, particularly in addressing both token- and sequence-level confidence and output consistency.
@@
  • We present a family of Confidence and Consistency-based Approaches (CoCoA) to UQ, offering a new way to merge information- and consistency-based measures for uncertainty quantification in LLMs.
@@
  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and translation. Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
Figure 2: Illustration of the method: the LLM generates a response, evaluates the similarity to alternatives, computes the uncertainty, and combines the uncertainty score with the similarity measures. High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.
@@ 2 Background
In this section, we introduce key concepts related to uncertainty quantification for LLMs, outline existing methods, discuss their limitations, and highlight the motivation for our approach.
First and foremost, it is important to establish the concept of an uncertainty function. Let $\mathbf{y}=f(\mathbf{x})$ denote the output of an LLM given an input sequence $\mathbf{x}$. The model defines a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$, from which outputs can be sampled. An uncertainty function $U$ is a mapping that quantifies the level of uncertainty $u$ associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as
$u=U(\mathbf{y}\mid\mathbf{x}).$ (1)
With this notion in mind, we now discuss existing methods for uncertainty quantification, outline their limitations, and describe how our approach addresses them.
2.1 Single-Sequence Information-Based Methods
Information-based methods rely on a single sample from the LLM and estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. One of the simplest techniques of this kind is Maximum Sequence Probability (MSP):
$u_{\text{MSP}}=-\log p\bigl{(}\mathbf{y}\mid\mathbf{x}\bigr{)}.$ (2)
Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix C.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
@@ 2.1 Single-Sequence Information-Based Methods
    First, LLMs provide us with the probability of a specific answer, even though the same meaning could often be expressed in multiple ways. Therefore, to obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
@@
    Second, answers of LLMs in general have very low probabilities due to the size of the vocabulary and the length of the generated sequences. Therefore, their estimates by LLM are not very reliable in general.
@@
    Third, these methods do not provide information about the flatness or the variability of the answer distribution.
2.2 Consistency-Based Methods
@@
The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$, where ${\mathbf{y}^{(i)}\sim p(\mathbf{y}\mid\mathbf{x})}$. Consistency-based uncertainty quantification methods rely only on the diversity of answers $\mathbf{y}^{(i)}$ sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require knowledge about the probability distribution of the tokens and can be applied in the black-box setting, when only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, given $M$ samples from the model, consistency-based methods compute a similarity matrix $G$, where each element $g_{ij}$ represents some form of similarity between the sampled outputs $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$:
$g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}\in[0,1].$ (3)
The value $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=1$ indicates the complete equivalence between $\mathbf{y}^{(i)}$ and $\mathbf{y}^{(j)}$, and $g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}=0$ indicates that there is no similarity.
Similarity could be computed in various ways. For example, Lexical Similarity (Fomicheva et al., 2020) is the surface form similarity, which calculates the overlap of words or phrases in the generations. More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024). For example, the Degree Matrix approach considers a similarity matrix $G$, which is computed using a model for Natural Language Inference (NLI), which predicts the probabilities of entailment, ${p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})$, and contradiction, ${p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$, between pairs of sentences $\mathbf{y}$ and $\mathbf{y}^{\prime}$. The similarity between two sequences is then defined as either $g_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})={p}_{\mathrm{entail}}(\mathbf{y},\mathbf{y}^{\prime})\quad\text{or}\quad g_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})=1-{p}_{\mathrm{contra}}(\mathbf{y},\mathbf{y}^{\prime})$. Lin et al. (2024) proposes to compute an averaged similarity matrix as $g_{ij}=\bigl{(}g\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}+g\bigl{(}\mathbf{y}^{(j)},\mathbf{y}^{(i)}\bigr{)}\bigr{)}/2$. The diagonal matrix $D$ is then defined with elements $D_{ii}=\sum_{j=1}^{M}g_{ij}$. The uncertainty from this soft clustering of meanings can then be defined as follows:
$U_{\mathrm{DegMat}}=1-\frac{\mathrm{trace}(D)}{M^{2}}.$ (4)
@@
The advantage of these techniques is that by generating multiple samples and analyzing their semantic similarity, they can obtain empirical probabilities for meanings instead of individual answers. The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
2.3 Information-Based Methods with Repeated Sampling
Averaging uncertainties.
The uncertainty scores can be aggregated using simple Monte Carlo averaging:
$u_{\text{MC}}=\frac{1}{M}\sum_{i=1}^{M}u_{i}.$ (5)
For the case when using the MSP uncertainty measure, i.e., when $u_{i}=-\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$, we obtain $u_{\text{MC}}=-\frac{1}{M}\sum_{i=1}^{M}\log p\bigl{(}\mathbf{y}^{(i)}\mid\mathbf{x}\bigr{)}$. The other notable example is the Monte Carlo Sequence Entropy (Kuhn et al., 2023).
While simple averaging represents a natural way to aggregate uncertainties, it has certain issues related to the nature of LLMs. First of all, in the vast majority of applications, an LLM-based system should produce a single output $\mathbf{y}_{*}$ for an input query. When we consider $u_{\text{MC}}$, we essentially perform averaging of uncertainties of different sequences, thus somewhat assessing the uncertainty related to the entire generative distribution $p(\mathbf{y}\mid\mathbf{x})$ for the input $\mathbf{x}$, but not for a particular generated sequence $\mathbf{y}_{*}$. This averaged uncertainty might not be adequate for this particular sequence and, remarkably, often performs worse than the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$, which is related solely to the output $\mathbf{y}_{*}$. Moreover, although intuitive, this naïve aggregation method assumes that all outputs contribute equally to the final uncertainty estimate, regardless of their semantic relationships. This can lead to inconsistencies when semantically equivalent outputs have varying uncertainty scores or when outputs with low similarity are treated as equally important.
Semantically weighted averaging.
The basic idea of aggregation approaches like Semantic Entropy (Kuhn et al., 2023) or SAR (Duan et al., 2024) is to perform a weighted averaging of output probabilities and give more weight to sequences semantically similar to the response shown to a user. All recently proposed techniques, such as SAR and Semantic Entropy, can be unified into a semantically-aware Generalized Monte Carlo uncertainty estimate, defined as
$u_{\text{GMCU}}=\frac{1}{M}\sum_{i=1}^{M}h\Biggl{(}\sum_{j=1}^{M}g_{ij}\,p_{j}\Biggr{)}.$ (6)
Here, the inner summation aggregates sequence probabilities $p_{j}$ weighted by their semantic similarity to the $i$-th output, and the outer summation averages these contributions across all samples. The function $h(\cdot)$ provides an additional layer of flexibility, transforming the reweighted uncertainty scores, making the method a generalized framework for uncertainty quantification. Existing methods, such as Semantic Entropy and SAR, can be considered as special cases of this more comprehensive approach, where the functions $h$ and $g$ are chosen appropriately.
Unfortunately, methods that fall under GMCU, while offering benefits, also inherit the aforementioned issues from both categories of methods:
  1. 1.
    The term $\sum_{j=1}^{M}g_{ij}\,p_{j}$ aims to average the probabilities of semantically similar sequences to obtain more robust estimate of the probability. However, due to the extreme instability of the LLM probabilities, as shown in Figure 1, the aggregated probabilities often perform worse than non-aggregated baselines.
  2. 2.
    The outer summation in (6), similarly to the case of simple Monte Carlo averaging (5), often fails to outperform the uncertainty $u_{*}=U(\mathbf{y}_{*}\mid\mathbf{x})$ of a single generated sequence $\mathbf{y}_{*}$.
Let us note that all uncertainty functions discussed so far have the following properties:
  • Non-Negativity. The uncertainty function produces nonnegative values, that is, $U(\mathbf{y})\geq 0$ for all $\mathbf{y}$.
  • Monotonicity. Higher values of the uncertainty function $U$ indicate higher uncertainty, i.e., if output $\mathbf{y}^{(1)}$ is considered more uncertain than output $\mathbf{y}^{(2)}$, then $U\bigl{(}\mathbf{y}^{(1)}\bigr{)}\geq U\bigl{(}\mathbf{y}^{(2)}\bigr{)}$.
These properties become important to perform the kind of synthesis of confidence and consistency we propose in the following section.

### v3

# Uncertainty Quantification for LLMs through Minimum Bayes Risk: Bridging Confidence and Consistency

@@ Abstract
Uncertainty quantification (UQ) methods for Large Language Models (LLMs) encompass a variety of approaches, with two major types being particularly prominent: information-based, which focus on model confidence expressed as token probabilities, and consistency-based, which assess the semantic relationship between multiple outputs generated using repeated sampling. Several recent methods have combined these two approaches to boost UQ performance. However, they sometimes fail to outperform much simpler baseline methods. Our work discusses the fundamental approach to constructing uncertainty measures that directly links uncertainty with the minimum Bayes risks achieved by LLM decoding. Building on these findings, we propose a novel approach to integrating model confidence with output consistency, resulting in a family of efficient and robust UQ methods. We evaluate our approach across various tasks such as question answering, abstractive summarization, and machine translation, demonstrating sizable improvements over state-of-the-art UQ approaches.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized natural language processing (NLP), enabling advances in information retrieval Zhu et al. (2023), question answering Kwiatkowski et al. (2019), machine translation Kocmi and Federmann (2023), and a broad range of other NLP applications. As these models become an integral part of our everyday life, ensuring the reliability of their outputs is crucial, especially in high-stakes scenarios where errors can have serious consequences. One way to address this challenge is through uncertainty quantification (UQ), which estimates the reliability of model outputs and enables abstention from generations that are likely to be incorrect.
UQ for LLMs is a rapidly advancing research area, with new UQ methods emerging each year. Most novel techniques are based on two fundamental approaches: (a) information-theoretic analysis or (b) assessment of output consistency.
@@
Information-theoretic methods quantify the confidence of a model by analyzing the probability distributions it induces for predictions (Malinin and Gales, 2021; Fomicheva et al., 2020). A key limitation of these methods is that they cannot account for semantic variability across multiple possible responses for the same input. Specifically, the model may generate answers with the same meaning but with very different assigned probabilities; see Figure 1. LLMs are trained to predict the next token in a sequence based on patterns observed in vast amounts of data, resulting in varying probabilities for semantically equivalent output sequences.
In contrast, consistency-based methods directly analyze the semantic relationship between the sampled outputs (Fomicheva et al., 2020; Lin et al., 2024), capturing uncertainty as objective variability of meaning among the sampled outputs.
Information-theoretic and consistency-based methods have complementary strengths. For this reason, recent state-of-the-art methods aimed to unify these approaches (Kuhn et al., 2023; Duan et al., 2024). We follow this direction and propose a way of quantifying risk as a combination of these basic measures. This results in a family of efficient and robust UQ techniques. Our approach combines the strengths of both information-based and consistency-based methods, providing a more comprehensive and accurate assessment of uncertainty.
Recently, it has been argued Wang and Holmes (2025); Daheim et al. (2025) that UQ in NLP tasks can be viewed through the lens of the Minimum Bayes Risk (MBR) framework. Following this, we formulate our methods as particular forms of risk functions under the MBR approach.
Our main contributions can be summarized as follows:
@@
  • We propose a new way of quantifying risk that combines information-theoretic and consistency-based measures, and derive a family of Confidence and Consistency-based Approaches (CoCoA) to uncertainty quantification in LLMs.
  • We show that these measures can be approximated by a learned function trained on unlabeled held-out set with a negligible loss of performance, eliminating the need for costly repeated sampling from the LLM.
@@
  • We evaluate our approaches across a variety of NLP tasks, including question answering, summarization, and machine translation. Our experiments demonstrate sizable improvements in the reliability and robustness of UQ compared to state-of-the-art methods.
@@ 2 Background
2.1 Language Model Decoding
LLMs define a probabilistic output distribution $p(\mathbf{y}\mid\mathbf{x})$ for the given input sequence $\mathbf{x}$. The standard ways to obtain a particular output $\mathbf{y}_{*}$ given $p(\mathbf{y}\mid\mathbf{x})$ include various variants of sampling $\mathbf{y}_{*}\sim p(\mathbf{y}\mid\mathbf{x})$ and greedy decoding, where
$\mathbf{y}_{*}=\arg\max_{\mathbf{y}}p(\mathbf{y}\mid\mathbf{x}).$ (1)
An alternative approach to decoding is Minimum Bayes Risk (MBR) decoding Kumar and Byrne (2004):
$\mathbf{y}_{*}=\arg\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x}),$ (2)
where $\mathcal{Y}$ is the set of candidate sequences and $R(\mathbf{y}\mid\mathbf{x})$ is a risk function:
$R(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\,\,r(\mathbf{y},\mathbf{y}^{\prime})$ (3)
with $r(\mathbf{y},\mathbf{y}^{\prime})$ being a pairwise loss function. The standard choice in the MBR literature is to take $r(\mathbf{y},\mathbf{y}^{\prime})\propto-s(\mathbf{y},\mathbf{y}^{\prime})$, where the so-called utility function $s(\mathbf{y},\mathbf{y}^{\prime})$ represents some notion of similarity between generated sequences $\mathbf{y}$ and $\mathbf{y}^{\prime}$. Below we discuss how various decoding strategies and corresponding risk functions lead to well-grounded definitions of uncertainty.
2.2 From Risk to Uncertainty
Generally, an uncertainty function $U$ is a mapping that quantifies the level of uncertainty associated with the output of a model $\mathbf{y}$, conditioned on the input sequence $\mathbf{x}$, which we denote as $U(\mathbf{y}\mid\mathbf{x})$.
Bayesian uncertainty measures are well known to be strongly connected with minimum risks of various kinds Xu and Raginsky (2022); Kotelevskii et al. (2025). Recently, this connection was revisited in the context of natural language generation Wang and Holmes (2025); Daheim et al. (2025), leading to the development of new MBR-based UQ methods. More precisely, one can consider the smallest achievable risk within $\mathcal{Y}$ or, equivalently, to maximum expected utility:
$U_{\text{MBR}}(\mathbf{y}_{*}\mid\mathbf{x})=\min_{\mathbf{y}\in\mathcal{Y}}R(\mathbf{y}\mid\mathbf{x})=R(\mathbf{y}_{*}\mid\mathbf{x}),$ (4)
where $\mathbf{y}_{*}$ is given by (2). Interestingly, MBR-based uncertainty (4) is applicable even to $\mathcal{Y}$ consisting of a single sequence $\mathbf{y}_{*}$.
Below, we discuss how various existing uncertainty measures can be seen as particular approximations of minimum Bayes risks for various choices of utility functions. We also derive new uncertainty measures based on the MBR framework.
Single Sequence Information-Based Methods. Information-based methods estimate the uncertainty of the generated sequence by aggregating the uncertainty scores of individual tokens. Within the framework of MBR decoding, consider $r(\mathbf{y},\mathbf{y}^{\prime})=\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}$. In this case, Bayes risk corresponds to the expected zero-one loss of decoding:
$R_{0/1}(\mathbf{y}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\mathbf{1}\{\mathbf{y}^{\prime}\neq\mathbf{y}\}=1-p(\mathbf{y}\mid\mathbf{x}).$ (5)
This leads to one of the simplest information-based uncertainty measures, Maximum Sequence Probability (MSP):
$U_{\text{MSP}}(\mathbf{y}_{*}\mid\mathbf{x})=1-p\bigl{(}\mathbf{y}_{*}\mid\mathbf{x}\bigr{)}.$ (6)
Although widely used as an uncertainty measure across various applications, including natural language generation, its connection to statistical risk has only recently been explored in the UQ literature Kotelevskii et al. (2022, 2025); Aichberger et al. (2024).
Several other measures fall into this category, including Perplexity and Mean Token Entropy (Fomicheva et al., 2020); see Appendix D.1 for details. While using only a single sample makes them computationally efficient, these techniques face three major challenges:
@@ 2.2 From Risk to Uncertainty
    LLMs provide us only the probability of a specific answer, even though the same meaning can often be conveyed in multiple ways. To obtain a proper probability for the meaning of an answer, we need to marginalize over its various possible rephrasings. However, this is not feasible if we generate only a single sample.
@@
    Reliability of such methods depends heavily on the underlying LLM being well-calibrated, which is a quality that is hard to define and harder to ensure.
@@
    These methods are point estimates that do not provide information about the shape of the output distribution.
@@
Semantic Consistency-Based Methods. The aforementioned issues lead to the development of consistency-based methods based on repetitive sampling from the LLM. Consider that we have sampled a set of outputs $\bigl{\{}\mathbf{y}^{(i)}\bigr{\}}_{i=1}^{M}$, where ${\mathbf{y}^{(i)}\sim p(\mathbf{y}\mid\mathbf{x})}$. Consistency-based UQ methods rely on the diversity of answers $\mathbf{y}^{(i)}$ sampled from the LLM. The idea is that if the model outputs similar answers for the same prompt over and over again, it is confident in its predictions; otherwise, it is uncertain. These techniques do not require a probability distribution estimated by LLM and can be applied in the black-box setting, when only the generated tokens are available. This case is quite common when LLMs are deployed as a service and are accessible through a limited API.
Formally, consistency-based method start from defining some similarity function $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}\in[0,1]$ between arbitrary generations $\mathbf{y}$ and $\mathbf{y}^{\prime}$. The value $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}=1$ indicates the complete equivalence between $\mathbf{y}$ and $\mathbf{y}^{\prime}$, and $s\bigl{(}\mathbf{y},\mathbf{y}^{\prime}\bigr{)}=0$ indicates that there is no similarity. Similarity could be computed in various ways. For instance, the Lexical Similarity method Fomicheva et al. (2020) relies on surface-form similarity, measuring the degree of word- or phrase-level overlap between generated texts. More advanced techniques propose various methods for taking into account the semantic similarity of the generated answers by hard or soft clustering (Lin et al., 2024).
Finally, given $M$ samples from the model, standard consistency-based methods compute a similarity matrix $S$, where $s_{ij}=s\bigl{(}\mathbf{y}^{(i)},\mathbf{y}^{(j)}\bigr{)}$. Then, various statistics of $S$ are computed in order to estimate uncertainty (Lin et al., 2024), see details in Appendix D.2. Alternatively, one can focus on a particular output $\mathbf{y}_{*}$ and compute uncertainty measures based on its similarities $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$ with other samples $\mathbf{y}^{(i)}$; see Lin et al. (2024) for more details.
Interestingly, it was recently shown Wang and Holmes (2025); Daheim et al. (2025) that the latter approach has direct relation to MBR decoding. Building on the MBR framework, one can define the utility function as $r(\mathbf{y},\mathbf{y}^{\prime})=1-s(\mathbf{y},\mathbf{y}^{\prime})$ and derive the following consistency-based uncertainty score:
$U_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\mathbb{E}_{\mathbf{y}^{\prime}\sim p(\mathbf{y}\mid\mathbf{x})}\bigl{[}1-s(\mathbf{y}_{*},\mathbf{y}^{\prime})\bigr{]}.$ (7)
Then the Monte Carlo approximation of the minimum Bayes risk is given by:
$\widehat{U}_{\text{cons}}(\mathbf{y}_{*}\mid\mathbf{x})=\frac{1}{M}\sum\nolimits_{i=1}^{M}(1-s_{*i}),$ (8)
where $s_{*i}=s\bigl{(}\mathbf{y}_{*},\mathbf{y}^{(i)}\bigr{)}$. In our ablation study, we show that such an uncertainty measure reliably outperforms consistency-based measures that aggregate the pairwise similarities of all the samples (see Appendix C.2).
@@
A key strength of consistency-based techniques is that, by generating multiple samples and analyzing their semantic similarity, they can estimate empirical probabilities over meanings rather than individual answers. The main drawback is that they discard the useful information that comes from the probability distribution represented by the LLM, including estimates of probabilities of specific answers.
Figure 2: Illustration of the method: the LLM generates a response, evaluates the similarity to alternatives, computes confidence, and combines confidence with the similarity measure. High similarity to alternatives reduces the uncertainty, while low similarity keeps it high.